OpenStudy (anonymous):

The height of the Empire State Building is 318.00 meters. If a stone is dropped from the top of the building, what is the stone's velocity just before it strikes the ground?

OpenStudy (anonymous):

use principle of conservation of energy. gravitational potential energy at the top of the building = kinetic energy just before it strikes the ground => mgh = 1/2 mv^2 the only unknown is v.

OpenStudy (anonymous):

h = 1/2 a t^2

OpenStudy (anonymous):

find t

OpenStudy (anonymous):

and subs in the eq V = u + at

OpenStudy (anonymous):

u =0 becoz the stone is dropped

OpenStudy (anonymous):

@Vaidehi09 what is m and v?

OpenStudy (anonymous):

assume m to be the mass of the body. and v is its velocity just before it strikes the ground.

OpenStudy (dls):

lol don't confuse that guy

OpenStudy (anonymous):

@DLS lol i know right!

OpenStudy (anonymous):

guys...using conservation is the easiest approach here!

OpenStudy (dls):

you don't have the required data to apply convo

OpenStudy (anonymous):

after simplification, u'll have gh = 1/2 v^2 u know g. u know h. find v. who says we don't have the req data?

OpenStudy (dls):

sorry i didn't thought that mass would eliminate

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

You can use energy or kinematics to solve the problem With energy is mgh=mv^2/2 where m-mass of the body ,v-velocity ,h is the height m>0 then you can eliminate it then v^2=2gh v= \[\sqrt{2gh}\]

OpenStudy (anonymous):

square root (2gh) so 2 x 318 x 9.81 =6239.16. Take square root gives 78.988 m/s. The derivation is from 1/2 mv^2 = mgh so mv^2 = 2mgh then v^2 = 2gh so v = SqRoot (2gh)